please Help!!!!!!!!!!!!!
a=bh/2 trying to find missing base
a= 81, h=9

Answer by stripepanther1989(2)

replace the factors with what you know.
A=bh(1/2)
81=b(9)(1/2)
divide both sides by nine
9=b(1/2)
multiply both sides by 2 two get rid of hte 1/2
18=base